One side of a square is decreased by three and the other side by 5 resulting in an area decreased by 185. What are the new dimensions?
Answer by thejackal(72)

First what do we know? let the side of the square be x
thus, initially the area of the square was x^2
the area of the new rectangle = (x-3)(x-5) = x^2 - 185
solve for x:
x^2 - 3x - 5x + 15 = x^2 - 185
x^2 - x^2 - 8x + 15 = -185
185 + 15 = 8x
200 = 8x
25 = x
the new dimensions are thus (25-3) = 22, and (25-5) = 20
check this by
initially, area was 25 * 25 = 625 so (25-3)(25-5)+185 should be = 625